OpenStudy (anonymous):

PLEASE HELP: MEDAL AND FAN Solve the triangle. A = 51°, b = 11, c = 7

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

can you help? @swamijatin2

OpenStudy (swamijatin2):

what we have to find perimeter or area

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

lol wrong thing sorry

OpenStudy (anonymous):

I need a (side a); C(angle c); and B(angle b)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

ok

OpenStudy (swamijatin2):

or well that doesnt have to be a right triangle but

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

how do i do this

OpenStudy (anonymous):

draw a triangle that looks reasonably like this one then use the law of cosines

OpenStudy (anonymous):

i have a triangle but how do i use the law of cosines (i know what it is just not how to do it)

OpenStudy (anonymous):

\[a^2=b^2+c^2-2bc\cos(A)\] \[a^2=11^2+7^2-2\times 11\times 7\cos(51^\circ)\] and a calculator

OpenStudy (swamijatin2):

To solve the traingle we need to find remaining side a and angles B and C. Use law of cosines a2 = b2 + c2 - 2bc cosA a2 = (11)2 + (7)2 - 2(11)(7) cos(51) a2 = 121 + 49 - 154 cos(51) a2 = 170 - 154(0.629) a2 = 170 - 96.866 a2 = 73.134 a = 8.55

OpenStudy (anonymous):

i got that ok so how about the and C

OpenStudy (anonymous):

B and C angles

OpenStudy (swamijatin2):

Now use law of sines sinA/a = sinB/b sin(51)/ 8.55 = sin(B)/11 sin(B) = 11[sin(51)/ 8.55] sin(B) = 11[0.777/8.55] B = sin-1(0.999) B = 87.43°

OpenStudy (anonymous):

ok and C?

OpenStudy (swamijatin2):

Sum of angles in triangle = 180° A + B + C = 180 C = 180 - (A + B) C = 180 - (51 + 87.43) C = 180 - 138.43 C = 41.57°

OpenStudy (anonymous):

because none of those are my answers

OpenStudy (swamijatin2):

then you are incorrect anywhere

OpenStudy (anonymous):

what?

OpenStudy (swamijatin2):

did u get the answers

OpenStudy (anonymous):

8.5 is a which is only for one answer choice where B = 85.9 and C = 39.1

OpenStudy (anonymous):

a ≈ 12.8, C ≈ 39.1, B ≈ 89.9 a ≈ 8.5, C ≈ 39.1, B ≈ 89.9 No triangles possible a ≈ 12.8, C ≈ 43.1, B ≈ 85.9

OpenStudy (swamijatin2):

you r correct but according to question we have to do

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

please finish helping me @swamijatin2

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

we have done a lot but it doesn't match my answer choices

OpenStudy (phi):

can you re-post the question and the answer you got? It would take a long time to read through all of the posts up above.

OpenStudy (anonymous):

sure hold on

OpenStudy (anonymous):

Solve the triangle. A = 51°, b = 11, c = 7

OpenStudy (anonymous):

Work From above: a2 = b2 + c2 - 2bc cosA a2 = (11)2 + (7)2 - 2(11)(7) cos(51) a2 = 121 + 49 - 154 cos(51) a2 = 170 - 154(0.629) a2 = 170 - 96.866 a2 = 73.134 a = 8.55 Now use law of sines sinA/a = sinB/b sin(51)/ 8.55 = sin(B)/11 sin(B) = 11[sin(51)/ 8.55] sin(B) = 11[0.777/8.55] B = sin-1(0.999) B = 87.43° Sum of angles in triangle = 180° A + B + C = 180 C = 180 - (A + B) C = 180 - (51 + 87.43) C = 180 - 138.43 C = 41.57°

OpenStudy (anonymous):

My answer choices are: a ≈ 12.8, C ≈ 39.1, B ≈ 89.9 a ≈ 8.5, C ≈ 39.1, B ≈ 89.9 No triangles possible a ≈ 12.8, C ≈ 43.1, B ≈ 85.9

OpenStudy (anonymous):

so... what do we do?

OpenStudy (phi):

side a is correct. a= 8.55 redo your work for angle B

OpenStudy (phi):

the problem is when you do the work for angle B, you need lots of decimals of accuracy. for side a, don't round it to 8.55. use all the digits your calculator has.

OpenStudy (anonymous):

ok

OpenStudy (phi):

it might be more accurate to first find angle C sin C = (7*sin51)/8.548956649

OpenStudy (anonymous):

what do you get for angle B i got the same answer

OpenStudy (anonymous):

I got 0.63633750 Radians for C?

OpenStudy (phi):

sin C = 0.63633750 C= asin( 0.63633750 ) = 39.519 deg

OpenStudy (phi):

summarizing a= 8.55, b= 11, c= 7 A= 51, B= 89.5, C= 39.5 the best match to your choices is choice B

OpenStudy (phi):

if we use a=8.6 (round from 8.55) we would get C= 39.2 and B= 89.8 which is even closer to choice B.

OpenStudy (phi):

because they use ≈ which means approximately it is reasonable to pick choice B. And the other choices are not "in the ball park"